# LICO: Large Language Models for In-Context Molecular Optimization

### v1

@@ 5.1 PMO Benchmark
The second best-performing method in this benchmark is GP BO, a method very similar to LICO, where the only difference between the two is the surrogate model. This indicates the superiority of LICO compared to GP, a popular surrogate model for black-box optimization. To verify this, we compare the predictive performance of LICO and GP on several objective functions. We do this by first labeling the ZINC unlabeled dataset with the objective functions and randomly choosing a subset of the labeled data points for evaluation. For each task, we vary the number of examples given to each method from $32$ to $512$, and evaluate their performance on $128$ held-out data points. We use negative log-likelihood, mean squared error, and root mean squared calibration error as the evaluation metrics. Figure 2 compares the predictive performance of LICO and GP in $3$ objective functions, median1, ranolazine_mpo, and troglitazone_rediscovery. The figure shows that the optimization performance of the method closely aligns with the predictive performance of the surrogate model. In median1 and ranolazine_mpo where LICO outperforms GP in terms of optimization score, the model also achieves lower negative log-likelihood, mean squared error, and calibration error. Similarly, LICO has worse predictive performance in troglitazone_rediscovery where it underperforms GP. This verifies our hypothesis and proves the effectiveness of LICO for surrogate modeling.
@@
Figure 2: The predictive performance of LICO and GP on $3$ objective functions in PMO with different metrics and varying numbers of observations.
@@
Table 2: Performance of LICO on $5$ tasks with different language instructions.
@@
Table 3: Performance of LICO on $5$ tasks with different ratios of synthetic data.
@@
LICO Synthetic $0.547\pm 0.080$ $0.498\pm 0.026$ $0.404\pm 0.103$ $0.585\pm 0.003$ $0.902\pm 0.012$ $2.936$
We also note that there is a discrepancy between Table 1 and the results reported in PMO. This is due to several reasons. First, we use a smaller optimization budget of $1000$ queries compared to $10000$ in PMO. We believe $1000$ is a more reasonable budget while still allowing optimization methods to achieve meaningful performances. Moreover, we found the GP BO implementation in PMO to be suboptimal, specifically in how it generated the candidate pool. The original implementation applied crossover and mutation to a mix of the best individuals and randomly selected individuals from the last iteration to generate the candidate pool for the current iteration. However, we found that only using the best individuals from the last iteration results in a much better performance. By implementing this change, GP BO becomes a much stronger baseline than Graph GA and REINVENT.
@@ 5.2.2 LICO with different synthetic ratios
We investigate the importance of training LICO on both intrinsic and synthetic data. To do this, we gradually increase the ratio of synthetic functions in the training data from 00 (intrinsic-only) to $1$ (synthetic-only), and compare the performance of LICO across different ratios. Table 3 shows that LICO with semi-synthetic training performs the best, outperforming both intrinsic-only and synthetic-only data. Training with synthetic data only performs the worst, which is expected when synthetic functions generated by a GP do not include any domain knowledge that is encoded by the intrinsic functions. In other words, synthetic data alone provides too little relevant signal for the model to generalize to unseen downstream objectives. Training with intrinsic functions only, on the other hand, results in quite good performances on most tasks. However, in tasks like albuterol_similarity, semi-synthetic training outperforms this baseline by a large margin. We hypothesize that the underlying objective in albuterol_similarity is far from the intrinsic functions used to train LICO, leading to poor generalization. Finally, training with small ($0.1$) to moderate ($0.5$) ratios of synthetic data achieves similarly good performance.
@@ 5.2.3 Randomly initialized vs Pretrained LLMs
To understand the importance of using a pretrained LLM, we compare LICO with an autoregressive transformer model of the same size (7B). The transformer architecture is the same as in [15], and we train this model to perform in-context learning on the semi-synthetic data from scratch. Table 4 shows the comparison. The scratch model performs much worse than LICO with a pretrained LLM on all tasks despite sharing the same number of parameters. This highlights the importance of the pattern-matching capabilities that LLMs like Llama-2 acquire through extensive language pretraining.
@@ 5.2.4 LICO with different LLM sizes
Previous works have shown the favorable scaling laws of Large Language Models where larger models consistently perform better on downstream tasks [24]. In this section, we investigate the scaling properties of LLMs but in the context of black-box optimization. Specifically, we compare $4$ different base LLMs with different sizes – Qwen-1.5 1.8B and 4B [4], Phi-2 2.7B [20], and Llama-2 7B [45]. We use the same language instructions for all models. We evaluate each model on the first $8$ tasks in Table 1 and average the results across $5$ random seeds. We report the sum of performance across $8$ tasks.
@@ 6 Conclusion and Future Work
We develop LICO, a new method that leverages pretrained Large Language Models for black-box optimization. LICO extends existing LLMs to non-language domains with separate embedding and prediction layers. To enable efficient generalization to various optimization tasks, we train LICO on a diverse set of semi-synthetic functions for few-shot predictions. LICO achieves state-of-the-art performance on PMO, a challenging molecular optimization benchmark with over 20 objective functions. Ablation analyses highlight the importance of incorporating language instruction to guide in-context learning and semi-synthetic training for better generalization.
One limitation of our method is the assumption of an accessible set of intrinsic functions. While this is true for molecular optimization, it may not apply to other scientific domains. In such cases, a better synthetic data generation process incorporating domain knowledge is needed to aid generalization.
Future directions include evaluating LICO in other domains to test its applicability and generality, exploring other prompts that better exploit the capabilities of pretrained LLMs, and using LLMs for other aspects of optimization, such as candidate suggestion or exploration.
Acknowledgements
This work is supported by grants from Amazon, Cisco, Microsoft, and Samsung.
@@ A.2 Training details
The $x$ embedding layer, $y$ embedding layer, and prediction layer in LICO are MLPs with a hidden dimension of $1024$. We train LICO for $20000$ steps with a batch size of $4$. For each data point in the batch, we randomly decide whether to sample an intrinsic or a synthetic function, with the probability of choosing synthetic functions being $0.1$. Each data point is a sequence of $(x,y)$ pairs with length $n\sim\mathcal{U}[64,800]$. If the function is an intrinsic function, we uniformly sample a property from Table 5, otherwise sample synthetic data following Equation (4).
We use Llama-2-7b [45] as the base LLM, and use LoRA [19] for parameter-efficient finetuning. We use a base learning rate of $45e-$ 4 with a linear warmup for $1000$ steps and a cosine decay for the remaining $19000$ steps. We use LoRA with a rank of $16$ and $\alpha$ scale of $16$.
@@ A.3 Black-box optimization hyperparameters
We use Algorithm 1 to optimize a black-box function with LICO. We initialize the observed dataset $\mathcal{D}_{\text{obs}}$ with a population of $34$ molecules sampled randomly from ZINC. At each iteration, we use the best $34$ candidates in $\mathcal{D}_{\text{obs}}$ to generate new candidates via crossover and mutation operations, with the mutation rate being $0.01$. The candidate pool size $C$ is $100$. We predict the mean $\mu_{i}$ and standard deviation $\sigma_{i}$ for each candidate $x_{i}$ in the pool using LICO. We employ a UCB acquisition function to compute the utility score $u_{i}=\mu_{i}+\beta\sigma_{i}$, which balances exploration and exploitation. Following [14], we set $\beta=10^{b}$, where $b\sim\mathcal{U}[-0.5,1.5]$. We then pick $k=15$ candidates with the highest utility scores. We evaluate each selected candidate $x_{j}$ using the black-box function $f$, and add the new data point $(x_{j},y_{j})$ to the observed dataset $\mathcal{D}_{\text{obs}}$. The process continues with the updated observed dataset, and stops when $|\mathcal{D}_{\text{obs}}|=1000.$
When predicting $\mu_{i},\sigma_{i}=f_{\theta}(x_{i},\mathcal{D}_{\text{obs}})$, we normalize all the $y^{\prime}s$ values in $\mathcal{D}_{\text{obs}}$ to have mean 00 and standard deviation $1$. This is to resemble the finetuning data distribution of LICO. We then denormalize $\mu_{i}$ and $\sigma_{i}$ to the original space.
@@ Appendix B Baseline details
is a transformer-based architecture for in-context learning. We refer to Nguyen and Grover [36] for more details about TNP. We train a TNP model with $16$ attention layers and $2048$ hidden dimensions. Other hyperparameters are the same as for LICO. After training, we use TNP for black-box optimization using Algorithm 1 with the same optimization hyperparameters but replace LICO with TNP.
@@
replaces the LICO surrogate model in Algorithm 1 with a Gaussian Process with a Tanimoto kernel. We optimize the Gaussian Process hyperparameters via maximum likelihood estimation on the initial population sampled from ZINC. Other optimization hyperparameters are the same as for LICO.
@@
is a model-free variant of Algorithm 1. Specifically, at each iteration, Graph GA generates $15$ candidates using the same crossover and mutation operations, and directly evaluates and adds them to $\mathcal{D}_{\text{obs}}$, since it does not employ a surrogate model.
@@
adopts a policy-based RL approach to finetune a pretrained RNN to generate SMILES strings with high returns. At each optimization iteration, we sample $16$ molecules from the finetuned RNN, evaluate these molecules using the black-box function $f$, and add the new data points to $\mathcal{D}_{\text{obs}}$. We refer to Gao et al. [14] for more details of the algorithm and other hyperparameters.
@@ C.1 Additional metrics
In addition to AUC Average Top-10, we measure the optimization performance of different methods on AUC Average Top-1 and AUC Average Top-100 for a more comprehensive comparison. Table 6 and 7 show AUC Average Top-1 and AUC Average Top-100 performances, respectively.
Table 6: The performance of LICO and the baselines on $21$ optimization tasks in PMO with AUC Average Top-1 metric. A higher score is better. We report the mean and stddev of scores averaged over $5$ random seeds. We use blue and violet to denote the best and second-best method for each task.
@@
valsartan_smarts ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.000\pm 0.000}}$ $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.000\pm 0.000}}$ $000.000\pm 0$ 0.000
@@
Table 7: The performance of LICO and the baselines on $21$ optimization tasks in PMO with AUC Average Top-100 metric. A higher score is better. We report the mean and stddev of scores averaged over $5$ random seeds. We use blue and violet to denote the best and second-best method for each task.
@@
scaffold_hop ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.462\pm 0.006}}$ $0.435\pm 0.008$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.462\pm 0.006}}$ $0.415\pm 0.009$ $0.440\pm 0.010$
@@
valsartan_smarts ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.000\pm 0.000}}$ $000.000\pm 0$ 0.000 $000.000\pm 0$ 0.000 ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.000\pm 0.000}}$ $000.000\pm 0$ 0.000

### v2

@@ 5.1 PMO Benchmark
The most closely related method to LICO is GP BO, where the only difference between the two is the surrogate model. This indicates the superiority of LICO compared to GP, a popular surrogate model for black-box optimization. To verify this, we compare the predictive performance of LICO and GP on several objective functions. We do this by first labeling the ZINC unlabeled dataset with the objective functions and randomly choosing a subset of the labeled data points for evaluation. For each task, we vary the number of examples given to each method from $32$ to $512$, and evaluate their performance on $128$ held-out data points. We use negative log-likelihood, mean squared error, and root mean squared calibration error as the evaluation metrics. Figure 2 compares the predictive performance of LICO and GP in $3$ objective functions, median1, ranolazine_mpo, and troglitazone_rediscovery. The figure shows that the optimization performance of the method closely aligns with the predictive performance of the surrogate model. In median1 and ranolazine_mpo where LICO outperforms GP in terms of optimization score, the model also achieves lower negative log-likelihood, mean squared error, and calibration error. Similarly, LICO has worse predictive performance in troglitazone_rediscovery where it underperforms GP. This verifies our hypothesis and proves the effectiveness of LICO for surrogate modeling.
@@
Figure 2: The predictive performance of LICO and GP on 33 objective functions in PMO with different metrics and varying numbers of observations.
@@
Table 3: Performance of LICO on 55 tasks with different language instructions.
@@
Table 4: Performance of LICO on 55 tasks with different ratios of synthetic data.
@@
LICO Synthetic $0.547\pm 0.080$ $98\pm 0.026 0.404$$04\pm 0.103$ $0.585\pm 0.003$ $0.902\pm 0.012$ $2.936$
@@ 5.2.2 LICO with different synthetic ratios
We investigate the importance of training LICO on both intrinsic and synthetic data. To do this, we gradually increase the ratio of synthetic functions in the training data from 0 (intrinsic-only) to $1$ (synthetic-only), and compare the performance of LICO across different ratios. Table 4 shows that LICO with semi-synthetic training performs the best, outperforming both intrinsic-only and synthetic-only data. Training with synthetic data only performs the worst, which is expected when synthetic functions generated by a GP do not include any domain knowledge that is encoded by the intrinsic functions. In other words, synthetic data alone provides too little relevant signal for the model to generalize to unseen downstream objectives. Training with intrinsic functions only, on the other hand, results in quite good performances on most tasks. However, in tasks like albuterol_similarity, semi-synthetic training outperforms this baseline by a large margin. We hypothesize that the underlying objective in albuterol_similarity is far from the intrinsic functions used to train LICO, leading to poor generalization. Finally, training with small ($0.1$) to moderate ($0.5$) ratios of synthetic data achieves similarly good performance.
@@ 5.2.3 Randomly initialized vs Pretrained LLMs
To understand the importance of using a pretrained LLM, we compare LICO with an autoregressive transformer model of the same size (7B). The transformer architecture is the same as in (Garg et al., 2022), and we train this model to perform in-context learning on the semi-synthetic data from scratch. Table 5 shows the comparison. The scratch model performs much worse than LICO with a pretrained LLM on all tasks despite sharing the same number of parameters. This highlights the importance of the pattern-matching capabilities that LLMs like Llama-2 acquire via extensive language pretraining.
@@ 5.2.4 LICO with different LLM sizes
Previous works have shown the favorable scaling laws of Large Language Models where larger models consistently perform better on downstream tasks (Kaplan et al., 2020). In this section, we investigate the scaling properties of LLMs but in the context of black-box optimization. Specifically, we compare $4$ different base LLMs with different sizes – Qwen-1.5 1.8B and 4B (Bai et al., 2023), Phi-2 2.7B (Javaheripi et al., 2023), and Llama-2 7B (Touvron et al., 2023b). We use the same language instructions for all models. We evaluate each model on the first $8$ tasks in Table 1 and average the results across $5$ random seeds. We report the sum of performance across $8$ tasks.
@@ 6 Conclusion and Future Work
We develop LICO, a new method that leverages pretrained Large Language Models for black-box optimization. LICO extends existing LLMs to non-language domains with separate embedding and prediction layers. To enable efficient generalization to various optimization tasks, we train LICO on a diverse set of semi-synthetic functions for few-shot predictions. LICO achieves state-of-the-art performance on PMO, a challenging molecular optimization benchmark with over 20 objective functions. Ablation analyses highlight the importance of incorporating language instruction to guide in-context learning and semi-synthetic training for better generalization. One limitation of our method is the assumption of an accessible set of intrinsic functions. While this is true for molecular optimization, it may not apply to other scientific domains. In such cases, a better synthetic data generation process incorporating domain knowledge is needed to aid generalization. Future directions include evaluating LICO in other domains to test its applicability and generality, exploring other prompts that better exploit the capabilities of pretrained LLMs, and using LLMs for other aspects of optimization, such as candidate suggestion or exploration.
@@ A.2 Training details
The $x$ embedding layer, $y$ embedding layer, and prediction layer in LICO are MLPs with a hidden dimension of $1024$. We train LICO for $20000$ steps with a batch size of $4$. For each data point in the batch, we randomly decide whether to sample an intrinsic or a synthetic function, with the probability of choosing synthetic functions being $0.1$. Each data point is a sequence of $(x,y)$ pairs with length $n\sim\mathcal{U}[64,800]$. If the function is an intrinsic function, we uniformly sample a property from Table 6, otherwise sample synthetic data following Equation equation 4.
We use Llama-2-7b (Touvron et al., 2023b) as the base LLM, and use LoRA (Hu et al., 2021) for parameter-efficient finetuning. We use a base learning rate of 5e45e-4 with a linear warmup for $1000$ steps and a cosine decay for the remaining $19000$ steps. We use LoRA with a rank of $16$ and $\alpha$ scale of $16$.
@@ A.3 Black-box optimization hyperparameters
We use Algorithm 1 to optimize a black-box function with LICO. We initialize the observed dataset $\mathcal{D}_{\text{obs}}$ with a population of $34$ molecules sampled randomly from ZINC. At each iteration, we use the best $34$ candidates in $\mathcal{D}_{\text{obs}}$ to generate new candidates via crossover and mutation operations, with the mutation rate being $0.01$. The candidate pool size $C$ is $100$. We predict the mean $\mu_{i}$ and standard deviation $\sigma_{i}$ for each candidate $x_{i}$ in the pool using LICO. We employ a UCB acquisition function to compute the utility score $u_{i}=\mu_{i}+\beta\sigma_{i}$, which balances exploration and exploitation. Following (Gao et al., 2022), we set $\beta=10^{b}$, where $b\sim\mathcal{U}[-0.5,1.5]$. We then pick $k=15$ candidates with the highest utility scores. We evaluate each selected candidate $x_{j}$ using the black-box function $f$, and add the new data point $(x_{j},y_{j})$ to the observed dataset $\mathcal{D}_{\text{obs}}$. The process continues with the updated observed dataset, and stops when $|\mathcal{D}_{\text{obs}}|=1000.$
When predicting $\mu_{i},\sigma_{i}=f_{\theta}(x_{i},\mathcal{D}_{\text{obs}})$, we normalize all the $y^{\prime}s$ values in $\mathcal{D}_{\text{obs}}$ to have mean 0 and standard deviation $1$. This is to resemble the finetuning data distribution of LICO. We then denormalize $\mu_{i}$ and $\sigma_{i}$ to the original space.
@@ Appendix B Baseline details
is a transformer-based architecture for in-context learning. We refer to Nguyen & Grover (2022) for more details about TNP. We train a TNP model with $16$ attention layers and $2048$ hidden dimensions. Other hyperparameters are the same as for LICO. After training, we use TNP for black-box optimization using Algorithm 1 with the same optimization hyperparameters but replace LICO with TNP.
@@
replaces the LICO surrogate model in Algorithm 1 with a Gaussian Process with a Tanimoto kernel. We optimize the Gaussian Process hyperparameters via maximum likelihood estimation on the initial population sampled from ZINC.
@@
is a model-free variant of Algorithm 1. Specifically, at each iteration, Graph GA generates a set of candidates using the same crossover and mutation operations, and directly evaluates and adds them to $\mathcal{D}_{\text{obs}}$, since it does not employ a surrogate model.
@@
adopts a policy-based RL approach to finetune a pretrained RNN to generate SMILES strings with high returns. At each optimization iteration, we sample a set of molecules from the finetuned RNN, evaluate these molecules using the black-box function $f$, and add the new data points to $\mathcal{D}_{\text{obs}}$. We refer to Gao et al. (2022) for more details of the algorithm and other hyperparameters.
B.1 Hyperparameter tuning for the baselines
To ensure the baselines achieve the best possible performance for the new PMO-1K benchmark, we performed extensive hyperparameter tuning for each baseline on the first $5$ tasks with grid search, and used the optimal hyperparameters for the rest of the tasks. Table 7 specifies the grid search for each method.
Table 7: Grid search and optimal hyperparameters for the baseline methods.
Method Grid Search Optimal Hyperparameters
GP BO population_size $\in$ {50, 100, 150, ..., 350}
offspring_size $\in$ {50, 100, 150}
kept_offspring_size $\in$ {5, 10, ..., 50}
population_size = 50
offspring_size = 100
kept_offspring_size = 15
Graph GA population_size $\in$ {50, 100, 150, ..., 350}
offspring_size $\in$ {50, 100, 150}
population_size = 50
offspring_size = 100
REINVENT batch_size $\in$ {4, 8, 16, 32, 64}
experience_replay $\in$ {4, 8, 16, 24, 32}
batch_size = 16
experience_replay = 24
Genetic GFN learning_rate $\in$ {0.0001, 0.0005}
batch_size $\in$ {4, 8, 16, 32, 64}
num_keep $\in$ {128, 256, 512, 1024}
offspring_size $\in$ {2, 4, 8}
ga_generations $\in$ {1, 2}
learning_rate = 0.0001
batch_size = 8
num_keep = 128
offspring_size = 2
ga_generations = 1
Augmented Memory batch_size $\in$ {4, 8, 16, 32, 64}
replay_buffer_size $\in$ {50, 100, 150}
batch_size = 32
replay_buffer_size = 100
@@ C.1 Additional metrics
In addition to AUC Average Top-10, we measure the optimization performance of different methods on AUC Average Top-1 and AUC Average Top-100 for a more comprehensive comparison. Table 8 and 9 show AUC Average Top-1 and AUC Average Top-100 performances, respectively.
Table 8: The performance of LICO and the baselines on $21$ optimization tasks in PMO with AUC Average Top-1 metric. A higher score is better. We report the mean and stddev of scores averaged over 55 random seeds. We use blue and violet to denote the best and second-best method for each task.
@@
valsartan_smarts $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.000\pm 0.000}$} $\pm 0.000 0.000±0.000$$\pm 0.000$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.000\pm 0.000}$} $0.000\pm 0.000$
@@
Table 9: The performance of LICO and the baselines on $21$ optimization tasks in PMO with AUC Average Top-100 metric. A higher score is better. We report the mean and stddev of scores averaged over 55 random seeds. We use blue and violet to denote the best and second-best method for each task.
@@
scaffold_hop $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.462\pm 0.006}$} $0.435\pm 0.008$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.462\pm 0.006}$} $15\pm 0.009 0.440$$40\pm 0.010$
@@
valsartan_smarts $\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mathbf{0.000\pm 0.000}$} $\pm 0.000 0.000±0.000$$\pm 0.000$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.000\pm 0.000}$} $0.000\pm 0.000$
@@
C.2 LICO with LLMs trained on molecular corpora
One may wonder whether using an LLM finetuned on molecular corpora helps improve the performance of LICO. To answer this question, we compare the performance of LICO with two different base LLMs: T5-base (Raffel et al., 2020) and Nach0-base (Livne et al., 2024), which finetunes T5-base on molecule corpora. Due to time constraints, we did not perform optimization with these new models, but compared their predictive performance instead. Figure 4 summarizes the results. There are two interesting observations from this figure. First, Nach0 works significantly better than plain T5, confirming the hypothesis that proper finetuning of a language model on molecule data helps boost its in-context property prediction in LICO. Second, Llama-2 works better than Nach0. As we explained in the paper, because we perform in-context learning in the embedding space of the language model, we rely on the general pattern-matching capability of the model, i.e., the ability to extract the relationship between embeddings of x and y from examples. From this perspective, it is not surprising that Llama-2 works better than a T5-based model, since it is much larger and has been pretrained with a lot more data, leading to a superior pattern-matching capability. One more benefit of using general LLMs like Llama is that they are domain-agnostic, which means we can finetune them for other non-molecule domains as well.
Figure 4: Predictive performance of LICO with T5-base, Nach0, and Llama-2 as the backbones.
C.3 LICO vs GPT-4 for molecular property prediction
One may wonder if we can prompt state-of-the-art LLMs like GPT-4 to perform molecular property prediction. To investigate this, we conducted an experiment where we prompted GPT-4o to perform in-context property prediction in the text space. The following text box shows the prompt we used.
I will give you a list of molecules and their corresponding property values. Based on these examples, your task is to predict the property value of a new molecule. Please provide your answer as a single number placed inside a pair of parentheses without any other information. For example, if you think the property value of the new molecule is 0.5, you should write 0.5. Molecule: [m1], Property: [p1] Molecule: [m2], Property: [p2] Molecule: [mn], Property:
Figure 5: Predictive performance of LICO vs GPT-4o.
Figure 5 compares the predictive performance of LICO with GPT-4o on 3 tasks – median1, ranolazine_mpo, and troglitazone_rediscovery, similar to the paper. We vary the context length from 32 to 512, and for each context length average the mean squared error across 128 target molecules. The result shows that prompting GPT4-o directly in the text space performs poorly, while LICO works much better and its performance improves as we increase the context length.
C.4 LLM Embeddings for molecular optimization
This work explored using LLMs as the in-context backbone for surrogate modeling in molecular optimization. Another way to make use of a pretrained LLM is to use the embeddings of an LLM as feature vectors for molecules and train a regressor on top of the embedding space. This section explores this idea with two regression models: a 4-layer MLP and a Gaussian Process (GP), both using the last hidden layer of the frozen LLama-2-7B model as the LLM embeddings. For the MLP baseline, we train the model on the full data buffer for 10 epochs with a batch size of $128$ and a learning rate of 1e41e-4, for every 150 molecules collected (every time after each new set of molecules is selected and labeled by the oracle). For the GP baseline, we keep the same hyperparameters as the GP BO method that uses fingerprint features. We compare LICO with these two baselines in the PMO benchmark.
Table 10: Comparison of LICO and other baselines with LLM embeddings on $23$ optimization tasks in PMO. A higher score is better. We report the mean and stddev of scores averaged over random seeds. We bold the best method for each task.
Task LICO MLP w/ LLM Embedding GP BO w/ LLM Embedding
albuterol_similarity $\mathbb{0.885\pm 0.019}$ $0.740\pm 0.149$ $0.868\pm 0.175$
amlodipine_mpo $\mathbb{0.679\pm 0.027}$ $0.621\pm 0.058$ $0.558\pm 0.025$
celecoxib_rediscovery $\mathbb{0.664\pm 0.122}$ $0.549\pm 0.113$ $0.633\pm 0.187$
deco_hop $\mathbb{0.619\pm 0.015}$ $0.594\pm 0.006$ $0.611\pm 0.007$
drd2 $0.928\pm 0.018$ $\mathbb{0.931\pm 0.194}$ $0.852\pm 0.263$
fexofenadine_mpo $\mathbb{0.772\pm 0.023}$ $0.734\pm 0.045$ $0.709\pm 0.053$
gsk3b $\mathbb{0.876\pm 0.045}$ $0.793\pm 0.034$ $0.845\pm 0.067$
isomers_c7h8n2o2 $0.939\pm 0.022$ $\mathbb{0.941\pm 0.116}$ $0.908\pm 0.170$
isomers_c9h10n2o2pf2cl $0.819\pm 0.039$ $\mathbb{0.835\pm 0.111}$ $0.739\pm 0.119$
jnk3 $\mathbb{0.731\pm 0.037}$ $0.725\pm 0.021$ $0.728\pm 0.029$
median1 $0.291\pm 0.016$ $0.282\pm 0.039$ $\mathbb{0.306\pm 0.044}$
median2 $0.280\pm 0.019$ $0.223\pm 0.026$ $0.280\pm 0.037$
mestranol_similarity $0.614\pm 0.064$ $\mathbb{0.770\pm 0.120}$ $0.612\pm 0.150$
osimertinib_mpo $\mathbb{0.820\pm 0.012}$ $0.809\pm 0.011$ $0.785\pm 0.012$
perindopril_mpo $\mathbb{0.557\pm 0.028}$ $0.497\pm 0.024$ $0.484\pm 0.034$
qed $0.936\pm 0.001$ $\mathbb{0.946\pm 0.002}$ $0.946\pm 0.002$
ranolazine_mpo $\mathbb{0.774\pm 0.008}$ $0.736\pm 0.085$ $0.715\pm 0.109$
scaffold_hop $\mathbb{0.547\pm 0.026}$ $0.478\pm 0.012$ $0.506\pm 0.010$
sitagliptin_mpo $\mathbb{0.567\pm 0.034}$ $0.544\pm 0.116$ $0.429\pm 0.129$
thiothixene_rediscovery $0.514\pm 0.037$ $0.430\pm 0.083$ $\mathbb{0.521\pm 0.125}$
troglitazone_rediscovery $\mathbb{0.380\pm 0.026}$ $0.299\pm 0.040$ $0.370\pm 0.083$
valsartan_smarts $\pm 0.000 0.000±0.000$$\pm 0.000$ $0.000\pm 0.000$
zaleplon_mpo $\mathbb{0.515\pm 0.017}$ $0.500\pm 0.037$ $0.483\pm 0.043$
Sum of scores ($\uparrow$) $\mathbb{14.708}$ $13.975$ $13.887$
Table 10 shows the superior performance of LICO against the two baselines, achieving the best performance in 14/23 tasks in PMO. In addition to the stronger empirical performance, a significant advantage of LICO is the ability to generalize to any objective function via in-context learning without finetuning.